OpenStudy (anonymous):

FAN AND MEDAL!! Given the parent function f(x) = 2x, which graph shows f(x) + 1? a) exponential function going through point 0, 2 and ending up on the right b)exponential function going through point 0, 0 and ending down on the right c)exponential function going through point 0, 0 and ending up on the right d)exponential function going through point 0, 1.5 and ending up on the right

OpenStudy (anonymous):

would it be a?

OpenStudy (mathmale):

It's important to express your functions correctly. At first I'd thought your 2x meant "2 times x." Please choose one or more of the following ways to express this exponential functin: f(x) = 2^x or \[f(x)=2^x\] or 

OpenStudy (anonymous):

the one you drew

OpenStudy (mathmale):

You'll need to draw the parent function, first. Then we can discuss what happens when you add 1 to f(x). The parent function is \[f(x)=2^x\]

OpenStudy (mathmale):

How would you graph this? Do y ou know the basic shape of exponential functions? this is a very important class of functions.

OpenStudy (anonymous):

curved right?

OpenStudy (anonymous):